OpenStudy (anonymous):

PLEASE HELP!!!! How do I write -8^(4/5), as a rational exponent?

OpenStudy (anonymous):

it is a rational exponent

OpenStudy (anonymous):

I know, but I need it to be further simplified, if it can.

OpenStudy (anonymous):

Actually, I got it out of a radical - I'm going from radical from, to rational exponent form.

OpenStudy (anonymous):

actually i am wrong, you would write \[-\sqrt[5]{8^4}\] you cannot do anything else

OpenStudy (anonymous):

Oh, okay - so -8^(4/5) cannot be simplified any more?

OpenStudy (anonymous):

I thought, where there is a negative 8, I would make it into a fraction.

OpenStudy (anonymous):

hold on \[\sqrt[n]{x^m}\] is radical form \[x^{\frac{m}{n}}\] is exponential form you have it in exponential form you do not know what the fifth root of 8 is (the cube root is 2) so you cannot simplify this further. your exponent is \(\frac{4}{5}\) and you cannot reduce this fraction, so you are done

OpenStudy (anonymous):

OH, I see! Okay, that makes sense. Thank you very much! :)

OpenStudy (anonymous):

also don't be confused about the minus sign. if it was in the EXPONENT that would mean take the reciprocal for example \[8^{-\frac{1}{3}}=\frac{1}{2}\] but the minus sign out front just means make it negative

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

OH, I see! That makes a lot of sense. I always get confused over whether or not it's the exponent or the base that turns it into a fraction.

OpenStudy (anonymous):

only if the minus sign is in the exponent do you take the reciprocal actually you already know this because i am sure you do not believe that \(-2=\frac{1}{2}\)

OpenStudy (anonymous):

Ha, ha - yes, that's true! :) Thanks!